# NSSI-Net: A Multi-Concept GAN for Non-Suicidal Self-Injury Detection Using High-Dimensional EEG in a Semi-Supervised Framework

### v2

@@ I Introduction
Electroencephalography (EEG) provides a direct method for reflecting neuronal dynamics originating in the central nervous system[8], making it highly valuable in detecting abnormal brain activities. EEG has been widely used across various disease fields due to its non-invasive nature, high temporal resolution, and sensitivity to brain dynamics. For example, it plays a crucial role in epilepsy research [9], Alzheimer’s Disease (AD) studies [10], depression diagnosis and monitoring [11], autism spectrum disorder (ASD) analysis [12], and schizophrenia research [13]. These applications underscore the versatility and importance of EEG in understanding and managing neurological and psychiatric conditions. In addition, the extraction of spatiotemporal features from EEG signals is crucial in revealing brain activity patterns related to diseases. For example, in seizure detection, EEGWaveNet developed a multiscale convolutional neural network (CNN) to extract spatiotemporal features [14], leading to significant improvements in detection accuracy. Liu et al. proposed a three-dimensional convolutional attention neural network (3DCANN) for EEG emotion recognition, integrating spatio-temporal feature extraction with an EEG channel attention module to effectively capture dynamic inter-channel relationships over time [15]. Zhang et al. introduced two deep learning-based frameworks that utilize spatio-temporal preserving representations of raw EEG streams to accurately identify human intentions. These frameworks combine convolutional and recurrent neural networks, effectively exploring spatial and temporal information either in a cascade or parallel configuration [16]. Gao et al. developed a novel EEG-based spatial-temporal convolutional neural network (ESTCNN) designed to detect driver fatigue by leveraging the spatial-temporal structure of multichannel EEG signals [17]. These studies highlight the effectiveness of spatiotemporal feature extraction in enhancing both the interpretability and performance of EEG-based disease studies.
@@ II-C Deep Learning for Feature Extraction from High-Dimensional EEG Data
Efficient feature extraction from high-dimensional EEG signals is essential for unraveling the complexities of brain functions and advancing accurate computational modeling. Traditional machine learning methods often struggle to capture the inherent and dynamic nature of EEG signals [24]. Deep learning methods offer the capability to automatically identify and characterize informative spatio-temporal features within EEG data, which not only enhance the precision of feature extraction but also allow for the dynamic adaptation to the non-linear and non-stationary characteristics of brain activity [27, 28, 29]. For example, a three-dimensional convolutional neural network (3D-CNN) was introduced, which processed EEG data as volumetric sequences and enhanced feature representation capabilities [30]. Considering the temporal dynamics in EEG series, a hierarchical spatial-temporal neural network was proposed to analyze the transition from regional to global brain activity, demonstrating the efficacy of deep learning in capturing detailed spatiotemporal information [31]. Besides of supervised learning, Liang et al. introduced a deep unsupervised autoencoder (AE) model (EEGFuseNet) designed to automatically characterize the interplay between spatial and temporal information in EEG data [32]. The latent features extracted by EEGFuseNet showed great potential for generalization across various applications. These works have underscored the transformative potential of deep learning in EEG research, which provide a more comprehensive, accurate and reliable understanding of brain activities.

### v3

@@ I Introduction
Electroencephalography (EEG) provides a direct method for reflecting neuronal dynamics originating in the central nervous system[8, 9], making it highly valuable in detecting abnormal brain activities. EEG has been widely used across various disease fields due to its non-invasive nature, high temporal resolution, and sensitivity to brain dynamics. For example, it plays a crucial role in epilepsy research [10], Alzheimer’s Disease (AD) studies [11], depression diagnosis and monitoring [12], autism spectrum disorder (ASD) analysis [13], and schizophrenia research [14]. These applications underscore the versatility and importance of EEG in understanding and managing neurological and psychiatric conditions. In addition, the extraction of spatiotemporal features from EEG signals is crucial in revealing brain activity patterns related to diseases. For example, in seizure detection, EEGWaveNet developed a multiscale convolutional neural network (CNN) to extract spatiotemporal features [15], leading to significant improvements in detection accuracy. Liu et al. proposed a three-dimensional convolutional attention neural network (3DCANN) for EEG emotion recognition, integrating spatio-temporal feature extraction with an EEG channel attention module to effectively capture dynamic inter-channel relationships over time [16]. Zhang et al. introduced two deep learning-based frameworks that utilize spatio-temporal preserving representations of raw EEG streams to accurately identify human intentions. These frameworks combine convolutional and recurrent neural networks, effectively exploring spatial and temporal information either in a cascade or parallel configuration [17]. Gao et al. developed a novel EEG-based spatial-temporal convolutional neural network (ESTCNN) designed to detect driver fatigue by leveraging the spatial-temporal structure of multichannel EEG signals [18]. These studies highlight the effectiveness of spatiotemporal feature extraction in enhancing both the interpretability and performance of EEG-based disease studies.
@@ II-C Deep Learning for Feature Extraction from High-Dimensional EEG Data
Efficient feature extraction from high-dimensional EEG signals is essential for unraveling the complexities of brain functions and advancing accurate computational modeling[28]. Traditional machine learning methods often struggle to capture the inherent and dynamic nature of EEG signals [25]. Deep learning methods offer the capability to automatically identify and characterize informative spatio-temporal features within EEG data[29], which not only enhance the precision of feature extraction but also allow for the dynamic adaptation to the non-linear and non-stationary characteristics of brain activity [30, 31, 32]. For example, a three-dimensional convolutional neural network (3D-CNN) was introduced, which processed EEG data as volumetric sequences and enhanced feature representation capabilities [33]. Considering the temporal dynamics in EEG series, a hierarchical spatial-temporal neural network was proposed to analyze the transition from regional to global brain activity, demonstrating the efficacy of deep learning in capturing detailed spatiotemporal information [34]. Besides of supervised learning, Liang et al. introduced a deep unsupervised autoencoder (AE) model (EEGFuseNet) designed to automatically characterize the interplay between spatial and temporal information in EEG data [35]. The latent features extracted by EEGFuseNet showed great potential for generalization across various applications. These works have underscored the transformative potential of deep learning in EEG research, which provide a more comprehensive, accurate and reliable understanding of brain activities.
@@ VI-B Evaluating Model Performance under Varying Labeled Source Data Ratios in Semi-Supervised Learning
Ten-fold cross-validation across subjects demonstrates the generalization ability of NSSI-Net. On the other hand, identifying the minimum number of training subjects remains important for computational efficiency. As shown in Fig. 5, performance improves significantly with around 10% labeled data (about 5 subjects) and continues to rise until about 75% (around 40 subjects). Incorporating 10%–25% of subjects (5 13 in this study) into the labeled training set offers an optimal balance between accuracy and computational cost.
@@ VI-E Topographic Analysis of Channel Contributions to NSSI Classification
Gender-specific analyses are conducted to explore potential neurophysiological factors affecting model performance, with the aim of enhancing scientific understanding rather than implying normative differences. Existing literature highlights biological and cognitive variations between males and females, particularly in emotional regulation and self-referential cognition[67, 68], which are key domains in the study of NSSI-related neural dynamics.